OpenStudy (anonymous):

the amount in a savings account increased from $200-$250 what was the percent of the increase?

OpenStudy (anonymous):

the actual increase is $50. now compare 50 to 200 by division: \[\frac{50}{200}=\frac{25}{100}=25\%\]

OpenStudy (anonymous):

If you use a simpel formula used in begining of economics at university level there is an equation \[V(1+r)^{n}\] where V is youre original amount of money, r is the interest and n is the period of time that it grows. By aplying it to your problem you get 200(1+r)^1 = 250 and by simply solving out the interest r you get 1+r = 250/200 = 1,25 => r = 0.25 or 25 %

OpenStudy (anonymous):

also a more basic way is [new - old]/old and then multiply by 100 to get a percent